it takes 20 mintes to do 8 math problems. if he has 28 problems how long will it take him
Answer by ptaylor(2198)

Let x=time it takes him to do 28 math problems
If it takes 20 minutes to do 8 math problems then he works at the rate of 8/20 of a math problem per minute (8/20th of a math problem per minute times 20 minutes equals 8 math problems). Soooo our equation to solve is:
(8/20)*x=28 multiply each side by 20
8x=560
x=70 minutes
CK
(8/20)*70=560/20=28
Hopes this helps---ptaylor